A proton and an alpha particle after accelerating through a potential difference Vand V2 volts respectively from rests have the same wavelength. The ratio of V1/V2 is 

Grade:11

3 Answers

Pooja
4 years ago
Relation of de Broglie wavelength and energy is
Lambda = h/p
E = p2/2m =  qV
therefore p = (2mqV)1/2
 So lambda * (mqV)1/2 = constant
 
Applying this for proton and alpha particle
mpqpV1 = maqaV2
V1/V2 = 4mp × 2qp / mp × qp
V1/V = 8
 
 
Khimraj
4 years ago
 
mpqpV1 = maqaV2
V1/V2 = 4mp × 2qp / mp × qp
V1/V = 8
…...................................................
Khimraj
4 years ago
 
mpqpV1 = maqaV2
V1/V2 = 4mp × 2qp / mp × qp
V1/V = 8
….............................................................